Shift Summary for Evening shift, Nov 4, 1999

Miguel Barrio (mbarrio@hep.uchicago.edu)
Thu, 4 Nov 1999 23:58:28 -0600

Shift Summary for Evening shift, Nov 4, 1999

On Shift: Ashkan, Miguelito

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

=========================================================================

Accounting of lost time:

No beam, sorry. Promises, promises, promises...

=========================================================================

Activities:

15:30 Running cosmics.

16:00 We get an announcement for beam around 18:00.

JasonL's futzing with the fruit, there's some sort of problem
with the kumquats.

17:00 (approx) JasonL stops the cosmic run and starts another one.

17:17 There's a message from MCR, beam is to be expected at around
20:00.

18:15 DN switches CO2 supply from dewar to bottle.

18:30 Analysis magnet back off (MCR checked for a power supply
problem since RickF had detected some bizarre smell).

19:15 MCR calls, they say beam should come our way soon.
We stop cosmic run and start muon run # 14905
(magnet's off and beam stop is in).

19:30 Upon starting triggers, KQ-3x acts up. Tried a coupla kfrends
to no avail so we resort to da ol' dirty method of cycling
da powa, which works. Turns out we also need to cycle the
laser crate repeater. All our trigger problems are solved now!

19:40 BobT notices there's no murmur window on the dart screen, so
we do a drt_fresh_start and wait for muon beam.

The problem with KQ-3y-6 not reading out seems to be solved, but
JasonL is skeptical and doesn't believe it won't show up again
(guess he's been betrayed by the fruits too many times to trust
them).

Meanwhile, KQ-4x is out of the readout, a new cable is needed.
There's the concern that this could affect the trigger rate as
the output of 4x is hard wired to the stiff track trigger.
Contacted Steve Bright and he says we could try just
disconnecting the cable that goes from 4x to the YTF. Will try
both configurations.

21:00 BobT calls MCR and requests an intensity of 9E12 (!!!).

Still no beam... :(

=========================================================================

Runs ended this shift: none

Tapes brought to FCC this shift: 0

=========================================================================

Duties for next shift:

**** IF BEAM COMES, PAGE BOBT.
Check with BobT if intensity requested is correct